# Towards Open Federated Learning Platforms: Survey and Vision from Technical and Legal Perspectives

### 18 Jan 2024 v2

@@ 4.3. How to Reuse Batch of Models
Once we have acquired a certain number of models that can contribute to the new target task, the next step is to reuse the knowledge of these pre-trained models, i.e., transfer their knowledge from source domain to the target domain (Pan and Yang, 2009). However, before deciding on how to reuse the model, it is important to ensure that the necessary legal rights and permissions have been obtained. This may involve reviewing the terms and conditions of the licenses under which the models were released or obtaining permission from the original creators or copyright holders. Therefore, in this section, we will not focus on the technical details of how to reuse models, which is already covered by many related surveys, such as Transfer Learning (Pan and Yang, 2009), Ensemble Leanring (Zhou, 2012), Domain Adaptation (Wang and Deng, 2018), Knowledge Distillation (Wang and Yoon, 2021), Deep Generative Models (Cao et al., 2022) and Model Fusion (Ji et al., 2021). Meanwhile, the specific model reuse technique or techniques used is at the user’s discretion, and the query-based FL platform we advocate is not bound or restricted to any particular model reuse algorithm. Innovatively, we study how to reuse batch of models, from the perspective of legal compliance.
@@
The machine learning community benefits from the openness of ideas and code, and many high-impact ML conferences and journals encourage authors to publish their source code and dataset to research platforms like Papers With Code  [^14] https://paperswithcode.com and Code Ocean  [^15] https://codeocean.com to increase exposure and facilitate reproducibility. To restrict the use of ML techniques for unethical purposes (i.e. Deepfakes (Mirsky and Lee, 2021)) and protect the IP of creators, models are typically published under a license agreed upon by the licensor. Here, we summary the granted rights, restrictions and enforcements of licenses for ML models posted on Hugging Face in Table 3. The following sections will provide a detailed survey of these licenses.
@@
Italian-Legal-BERT (Licari and Comandè, 2022)
@@
CodeGen (Nijkamp et al., 2023)/ A MIT-style license
@@ A MIT-style license
OpenRAIL ¿Responsible AI License template, w/o full text 22,947
ControlNet (Zhang et al., 2023a)
@@
Llama 2 (Touvron et al., 2023)
OpenRAIL++ ¿Same as CreativeML-OpenRAIL-M 1,433
Stable Diffusion v2 (Rombach et al., 2022)
BigScience-OpenRAIL-M ¿Same as BigScience-BLOOM-RAIL-1.0 659
@@
BLOOM (Scao et al., 2022)
BigCode-OpenRAIL-M ¿Same as BigScience-BLOOM-RAIL-1.0 320
StarCoder (Li et al., 2023a)
@@
SEER Vision Model (Goyal et al., 2022)
@@
*GFDL ¿Same as GPL, a free document license 30
@@
4.3.1. Model Licensing Forms
@@
As shown in Table 3, ML models are licensed in three main forms: as software (e.g. Apache, MIT, GPL), as a model (e.g. OpenRAIL), and as content/database (e.g. CC BY, PDDL). The reason for the mixed use of licenses is the ambiguity in the dependency relationship between the ML code, model, and data. Thinking in terms of software, ML models can be released with reproducable code and be considered as a component of software. So many Free and Open Source Software (FOSS) licenses (Rosen, 2005) are naturally deferred for licensing of models. The most popular license is Apache-2.0, which is a permissive FOSS license that allows the freedom to make derivative works. However, the model building process also relies on a massive amount of data (LeCun et al., 2015) that may be licensed under different licenses, which can lead to license conflicts. A real-world example is BERT (Devlin et al., 2019), which was published under the Apache-2.0 license but pre-trained on English Wikipedia documents that are licensed under CC BY-SA 3.0. This changing of license violates the requirement of the CC BY-SA 3.0, which states that any contribution must be distributed under the same license as the original work.
@@
Thinking in terms of content and database, some word embedding models, such as GloVe (Pennington et al., 2014), compute vector representations of words based on licensed open linguistic resources. These representations can be regarded as a translation of corpus and fall under the license of the original linguistic resources. A more complex scenario arises when the model is fine-tuned with other data that has a different license, for example, fine-tune RoBERTa (Liu et al., 2019) (licensed under MIT) with SQuAD2 (Rajpurkar et al., 2016) (licensed under CC BY 4.0). The resulting model can be interpreted as both derived works and combined works.
@@
Not only limited to protecting the IP and controlling the diffusion of ideas, but AI companies and researchers are also concerned about licensees using their models for unethical purposes (Jobin et al., 2019; Awad et al., 2018; Yuste et al., 2017), which is usually not restricted by traditional licenses designed for software and content. We can infer the concerns of unethical use of GPT-2 (Radford et al., 2019) from its modified MIT license granted by its inventors, which states, We don’t claim ownership of the content you create with GPT-2, so it is yours to do with as you please. We only ask that you use GPT-2 responsibly and clearly indicate your content was created using GPT-2. However, such a statement lacks legal enforcement, and users may avoid accountability by convincing themselves that despite their efforts to minimize harm, they could not predict the AI artifact they generated would be used for harmful purposes. On the other hand, the original licensing frameworks for software and content (e.g. MIT, CC BY) are not well suited to the data-driven ML. Many ML operations, such as training, fine-tuning, inference, and distillation, are not explicitly defined in traditional software and content licenses, leaving a potential legal loophole for licensees.
@@
To address the unique challenges and considerations surrounding the use and distribution of ML models, several specific licenses for ML models have been proposed. CreativeML OpenRAIL-M license, proposed by Responsible AI (Contractor et al., 2022), is the most popular model-specific license on Hugging Face and enables legally enforceable responsible use. By accepting this license, licensees must adhere to the use-based restrictions stated by the licensor, and these restrictions must also apply to derivative works. With a multitude of different model licenses available, it becomes a challenging and tedious work to reuse them in bulk. It is therefore imperative to establish guidelines for selecting the licenses for models and other related components that are ready for query-based FL.
4.3.2. License Choosing Preferences
@@
  • A model license ready for open FL paltforms should allow the modification, combination and redistribution of original works and any derived works; and
@@
  • Sublicensing right should be granted to lubricate the republication of derived works resulting from knowledge mining; note
@@
  • Some licenses enforce the source of the derived works to be disclosed and prohibit their commercial use, which hinders model selling (Chen et al., 2019); and
@@
  • Some licenses are copyleft (marked with * in Table. 3), which means the derivatives must be licensed under the same license or a compatible license, leading to potential license conflicts and proliferation (Gomulkiewicz, 2009); last
  • All granted rights are preferably irrevocable by the licensors (Reddy, 2009).
@@
Furthermore, it is important to consider the licensing of two other components when building and reusing models: data and algorithms, which may have entirely different license terms. Here, based on the comparisons between different licenses provided in Table 3, we present several strategies for selecting licenses in query-based FL to minimize conflicts.
Preferences for Datasets or Databases: CC0-1.0, PDDL, ODC-By ¿ CC BY, C-UDA ¿ LGPL-LR.
Our recommended licenses for training datasets and databases for query-based FL are CC0-1.0, PDDL, ODC-By and CC BY (the preferred version of CC BY is 4.0 due to the grant of Sui Generis Database Rights in Art.1c). CC0-1.0, PDDL and ODC-By are more permissive than CC BY since they do not require licensees to disclose any modifications made to the dataset or database. Additionally, CC0-1.0 and PDDL are public domain licenses and do not require the declaration of the original license. Although some of these licenses do not explicitly grant sublicensing rights, they provide an automatic licensing policy for downstream recipients.
C-UDA is an alternative license that grants sublicensing rights, but it includes additional usage restrictions that limit its application to computational use only, which indicates commercial use of data is not allowed. Nonetheless, C-UDA explicitly exempts reused results from any restrictions, which is highly favorable for our scenario of model mining. To avoid license proliferation, it is not recommended to use any data under copyleft licenses for building models, as the resulted models could be seen as remixing and making derivatives of the original datasets, leading to potential conflicts between licenses. Among them, LGPL-LR is an exception because it contains an exemption clause for work that uses the Linguistic Resource (Art.3), which is suitable for end-to-end training, fine-tuning, and embedding. But it is worth noting that the embedded representations may be considered translated straightforwardly into another language (Art.0), which falls within the scope of LGPL-LR license.
An example of license proliferation is LEGAL-BERT (Chalkidis et al., 2020), which was trained on data from the Case Law Access Project  [^16] https://case.law (licensed under CC BY-SA 4.0). This restricts LEGAL-BERT to the same license and prevents further model reusing on datasets or models licensed under incompatible copyleft licenses, such as LGPL-LR and GPL.
Preferences for Software: Apache-2.0, AFL-3.0, Artistic-2.0, ECL-2.0 ¿ MIT, BSD-3-Clause&-Clear, BSL-1.0, BSD-2-Clause, NCSA $\approx$ Ms-PL ¿ WTFPL-2.0, Unlicense, ISC, Zlib, PostgreSQL.
Our top recommended software licenses for training and reusing models are Apache, AFL, Artistic, and ECL. These permissive licenses allow modification and sublicensing, explicitly grant the use of patents and permit commercial use, and do not require the disclosure of the source code but only the stating of any changes made to the original work.
The next set of recommendations are MIT, BSD, BSL, and NCSA. These licenses do not explicitly grant patent rights but instead, do not require the stating of modifications made to the original work, thus avoiding the tedious task of tracking model reusing or incremental training procedures. Ms-PL offers two advantages simultaneously, but it is a weak copyleft license that requires the modified source code to also be licensed under Ms-PL, and the derivative object code to be compliant with a license compatible with Ms-PL. Note that FOSS licenses do not provide a clear definition for software-generated outputs such as models. It is unclear whether models are considered a portion of the software, and whether they are in source code form or object code form. This ambiguity makes it difficult to determine the applicable clauses for models.
Our latest recommended licenses include WTFPL, Unlicense, ISC, Zlib, and PostgreSQL. These licenses are very permissive and allow almost anything without restrictions. However, on the other side, these licenses also do not explicitly grant sublicensing rights and patent, which can lead to ambiguity in interpreting the license clauses. For the avoidance of doubt, copyleft licenses such as GPL, AGPL, LGPL, OSL, MPL, EPL, and EUPL are not recommended, despite the loophole that they do not have a specific definition for ML models. Although some of those copyleft licenses can be made compliant with others, we recommend isolating the software licenses from the resulting models to preserve the freedom to use the models further (e.g. close-source, relicense).
Preferences for Models: Apache-2.0, AFL-3.0, Artistic-2.0, ECL-2.0 ¿ OpenRAIL series.
There are two recommended choices for model licenses for query-based FL. The first is permissive FOSS licenses like Apache, AFL, Artistic, and ECL. The second is open model-specific licenses like OpenRAIL and its derivatives. As shown in Table 3, the main difference between the two choices is that OpenRAIL offers additional user behavioral restriction clauses and enforces these restrictions via a copyleft-style agreement. For example, CreativeML OpenRAIL-M license claims Therefore You cannot use the Model and the Derivatives of the Model for the specified restricted uses … You shall require all of Your users who use the Model or a Derivative of the Model to comply with the terms of this paragraph. The restricted uses include actions that could cause harm, provide medical advice, generate or disseminate verifiably false information, and more. So, the model owners may adopt these licenses for the purpose of responsible model use.
However, in practice, such discrimination of user behavior cannot completely guarantee that the models will not be misused, and may potentially compromise the openness of the models (Perens, 1999; Greenbaum, 2016). The user behavioral restrictions stated in licenses can be compared to manufacturers prohibiting the use of their laptops for hacking, and furthermore, the vendors can be held jointly and severally liable for any future violations, which is unreasonable. Therefore, including such statements in licenses may ultimately lead to the licensed materials becoming closed source. Additionally, to enable remote control for the responsible use of AI, CreativeML OpenRAIL-M includes the clause You shall undertake reasonable efforts to use the latest version of the Model, which requires licensees to keep up with the updates of the original work and may render their prior development efforts useless. Therefore, traditional permissive licenses, which follow worse-is-better design philosophy (Gabriel, 1991), are good choices for model licensing in query-based FL, as they promote openness and facilitate the sharing of publicly contributed models.
The remaining model licenses, OPT-175B and SEER, are proprietary licenses that allow licensees to use and reproduce the licensed models subject to certain restrictions. Given that their granted rights are revocable, we do not recommend using any content of works and derivatives under these licenses in query-based FL.
It is worth noting that the above discussion only deals with the licenses of inputs for open FL platforms, which aim to provide legal compliance and freedom of outputs as much as possible, but does not involve the copyright issue for the outputs. In fact, except for some public domain dedication licenses like CC0-1.0, PDDL, Unlicense, and WTFPL, most licenses only grant non-exclusive rights for use and distribution, and the original copyright and attribution are retained by the licensors. Whether the reused models are copyrightable is crucial for incentivizing model sharing and mining, so we will elaborate on this topic in the next section.
4.3.3. Copyright of Reused Models
Software and computer code are indisputably copyrightable, but what about computer-generated content such as distillation and ensembles of models? The copyrightable of a computer-generated work is controversial, which may depend on such as the level of creativity and originality and presence of at least minimal human creative effort at the time the work is produced (on New Technological Uses of Copyrighted Works, US). According to this definition, programmers who engage in model design and training meet the threshold requirements of copyrightability and own the copyright of the model. That is why all the licenses listed in Table 3 contain claims of copyright. But the debating point is whether the reused models also copyrightable? Unfortunately, there is no universal answer to this question as it can depend on the specific case and fact pattern. The crux is whether the efforts involved in reusing the model meet the minimum creative requirements for copyrightability. For example, if we simply stack two models end-to-end, it may not meet the threshold for copyrightability. However, if we improve a basis model using distilled knowledge from other domains, that would be more likely to meet the requirements for copyrightability. Except for copyrightability, the authorship of a reused model is also open to controversy, as it depends on whose original intellectual conceptions the work embodies, and joint authorship is also possible (Hedrick, 2019).
The determination of copyrightability and authorship of computer-generated content is an open issue that needs to be addressed through corresponding legislation (on New Technological Uses of Copyrighted Works, US; Hedrick, 2019; Margoni, 2018). European Parliament regarded that consideration must first be given to assessing patent law in the light of the development of AI  [^17] https://www.europarl.europa.eu/doceo/document/A-9-2020-0176_EN.html. The possible answers to the question of authorship of computer-generated models are model authors, model users, data owners, any combination of them, or no one (Hedrick, 2019). Licensors can also make efforts to clarify this issue by including relevant claims in their licenses. For example, the license of Stable Diffusion (Rombach et al., 2022) explicitly states that Licensor claims no rights in the Output You generate using the Model. Similarly, ChatGPT  [^18] https://openai.com/blog/chatgpt, even though it is a proprietary software of OpenAI company, its sharing & publication policy  [^19] https://openai.com/policies/sharing-publication-policy states The published content is attributed to your name or company. Therefore, we are free to use their generated content for model reusing and can claim the copyright of reused models. On the contrary, the licenses of OPT (Zhang et al., 2022c) and SEER (Goyal et al., 2022) do not grant any copyright for the data produced by the licensed software. Given that, we should avoid using their derivatives and generated content in query-based FL to prevent copyright infringement. Once we obtain the right to relicense the modification models, the choice of a new license depends on the application scenario of models. We further provide a flowchart in Fig. 4(a) to guide the license selection in the context of model query and model reusing.
@@ 4.3.3. Copyright of Reused Models
4.3.4. Batch Model Reuse Mechanisms
@@
Instead of summarizing the batch model reuse mechanisms from a technological and algorithmic aspect, we propose grouping these mechanisms based on the classification of their resulting outputs for ease of justifying license clauses. As shown in Fig. 4(b), there are four categories of batch model reuse mechanisms: Combination, Amalgamation, Distillation, and Generation, each resulting in different forms of outputs and may correspond to different regulations in licenses.
@@
Figure 4. (a) Flowchart for model licenses selection in the context of model query and model reusing. (b) Proposed taxonomy categorizing batch model reuse mechanisms based on the reused results.
[…]
@@
Combination (Zhou, 2012) is a straightforward way to reuse batch of models (aka base learners), in which multiple models jointly contribute to the output by combination strategies such as averaging, voting, learning (Jacobs et al., 1991; Wolpert, 1992). For regression estimates, averaging can improve the generalization by taking the mean of the outputs of all weak learners in a population. Additionally, the outputs of each learner can be weighted by extra parameters (Perrone and Cooper, 1995), which can be determined by stacking estimators (Wolpert, 1992), Bayes approach (Clarke, 2003) or backpropagation of gating networks (Jacobs et al., 1991). Voting is a workaround strategy for classification tasks and also applicable for stacking and gating. Both stacking and gating rely on an additional holdout or validation dataset for calculating extra parameters, marked as * in Fig. 4(b). The difference is that gating can adapt the weights of each model’s estimation based on the inputs, providing better generalizability performance of the combined model.
@@
There are many advantages of combination mechanisms from the perspective of FL. First, the input spaces of base models can be unaligned, which is ideal for the scenario of vertical FL (Wu et al., 2022a) where each client may have inconsistent features in their data. Secondly, especially for query-based FL, it can simultaneously support multiple types and heterogeneous models, which means that it does not rely on any prior assumptions of the models, such as whether they are DNNs or decision trees, released with raw weights (whitebox) or binary forms (blackbox). Thirdly, the tasks of models can be different if we pipeline the base models end-to-end, which is usually overlooked as a combination mechanism of models. Pipelining can fully leverage the transferability of models to solve previously unexplored ML problems. For instance, Gao et al. (Gao et al., 2022) proposed a zero-shot dense retrieval system named HyDE by pipelining a natural language generation (NLG) model (Ouyang et al., 2022) and a natural language understanding (NLU) model (Izacard et al., 2021). The generated content, which may lack factual grounding, from the NLG model is used as query embeddings to facilitate real document retrieval by the NLU model. Similarly, through query-based FL, we can query a vicarious NLG model for a novel scenario, such as ProGen (Madani et al., 2023) for protein sequences generation, and quickly adapt this system to proteomics. Not limited to that, we can query a batch of NLG models by a well-chosen filter condition and then combine models through averaging or gating to significantly expand the exploration space for knowledge discovery. Lastly, the combinated models have strong separation from each other, meaning that we can add or remove a batch of models without significant changes to the remaining ones. Meanwhile, combination mechanisms do not rely on the transparency of models and support blackbox sharing. Thus, the base model can establish loose connections with other models only through run scripts, providing revocability of such combination and circumvention of the restrictions of licenses.
@@
On the other hand, instead of being treated as a challenge for FL (Ma et al., 2022), the statistical heterogeneity and model heterogeneity nature of these crowdsourced models can actually enrich population diversity, which is crucial for creating a good ensemble (Maclin et al., 1995; Opitz and Shavlik, 1995). However, the storage consumption of combined models increases linearly with the number of base learners, which can strain the communication resources of a collaborative model training network. As a result, alternative approaches involve amalgamating or merging multiple models to create a new consensus model. In the following, we provide a summary of these methods.

### 29 Feb 2024 v3

@@ IV-C Legal Considerations in Batch Model Reuse
Once we have acquired a batch of models that can contribute to the new target task, the next step is to reuse the knowledge of these pre-trained models, i.e., transfer their knowledge from source domain to the target domain [80]. However, before deciding on how to reuse the model, it is important to ensure that the rights and permissions have been obtained. This may involve reviewing the terms and conditions of the licenses under which the models were released or obtaining permission from the original creators or copyright holders. Therefore, in this section, we will not focus on the technical details of how to reuse models, which is already covered by many related surveys, such as Transfer Learning [80], Ensemble Leanring [81], Domain Adaptation [82], Knowledge Distillation [83], Deep Generative Models [84] and Model Fusion [85]. Meanwhile, the specific model reuse techniques used are at the user’s discretion, and the query-based FL platform is not bounded or restricted to any particular reuse method. Innovatively, we study how to reuse batch of models, from the perspective of legal compliance.
@@
The machine learning community benefits from the openness of ideas and code, and many high-impact ML conferences and journals encourage authors to publish their source code and dataset to research platforms like Papers With Code and Code Ocean [^2] https://paperswithcode.com; https://codeocean.com to increase exposure and facilitate reproducibility. To restrict the use of ML techniques for unethical purposes (i.e., Deepfakes [86]) and protect the IP of creators, models are typically published under a license agreed upon by the licensor. Here, we summarized the granted rights, restrictions and enforcements of licenses for ML models posted on Hugging Face in TABLE II. The following sections will provide a detailed survey of these licenses.
@@
Italian-Legal-BERT [87]
@@
CodeGen [89]/ A MIT-style license
@@ A MIT-style license
OpenRAIL >Responsible AI License template, w/o full text 22,947
ControlNet [90]
@@
Llama 2 [92]
OpenRAIL++ >Same as CreativeML-OpenRAIL-M 1,433
Stable Diffusion v2 [91]
BigScience-OpenRAIL-M >Same as BigScience-BLOOM-RAIL-1.0 659
@@
BLOOM [73]
BigCode-OpenRAIL-M >Same as BigScience-BLOOM-RAIL-1.0 320
StarCoder [93]
@@
SEER Vision Model [95]
@@
*GFDL >Same as GPL, a free document license 30
@@
IV-C1 Model Licensing Forms
@@
As shown in TABLE II, ML models are licensed in three main forms: as software (e.g., Apache, MIT, GPL), as a model (e.g., OpenRAIL), and as content/database (e.g., CC BY, PDDL). The reason for the mixed use of licenses is the ambiguity in the dependency relationship between the ML code, model, and data. Thinking in terms of software, ML models can be released with reproducable code and considered as a component of software. So many Free and Open Source Software (FOSS) licenses [101] are naturally deferred for licensing of models. The most popular license is Apache-2.0, which is a permissive FOSS license that allows the freedom to make derivative works. However, the model building process also relies on a massive amount of data [102] that may be licensed under different licenses, which can lead to license conflicts. A real-world example is BERT [72], which was published under the Apache-2.0 license but pre-trained on English Wikipedia documents that are licensed under CC BY-SA 3.0. This changing of license violates the requirement of the CC BY-SA 3.0, which states that any contribution must be distributed under the same license as the original work.
@@
Thinking in terms of content and database, some word embedding models like GloVe [103], compute words representations based on licensed open linguistic resources. These representations can be regarded as a translation of corpus and fall under the license of the original linguistic resources. A more complex scenario arises when the model is fine-tuned with other data that has a different license, for example, fine-tune RoBERTa [104] (licensed under MIT) with SQuAD2 [97] (licensed under CC BY 4.0). The tuned model can be interpreted as both derived works and combined works.
@@
Not only limited to protecting the IP and controlling the diffusion of ideas, but AI companies and researchers are also concerned about licensees using their models for unethical purposes [105, 106, 107], which is usually not restricted by traditional licenses designed for software and content. We can infer the concerns of unethical use of GPT-2 [78] from its modified MIT license granted by its inventors, which states, We don’t claim ownership of the content you create with GPT-2, so it is yours to do with as you please. We only ask that you use GPT-2 responsibly and clearly indicate your content was created using GPT-2. However, such a statement lacks legal enforcement, and users may avoid accountability by convincing themselves that despite their efforts to minimize harm, they could not predict the AI artifact they generated would be used for harmful purposes. On the other hand, the original licensing frameworks for software and content (e.g., MIT, CC BY) are not well suited to the data-driven ML. Many ML operations, such as training, fine-tuning, inference, and distillation, are not explicitly defined in these license clauses, leaving a potential legal loophole for licensees.
@@
To address the unique challenges and considerations surrounding the use and distribution of ML models, several specific licenses for ML models have been proposed. CreativeML OpenRAIL-M license, proposed by Responsible AI [108], is the most popular model-specific license on Hugging Face and enables legally enforceable responsible use. By accepting this license, licensees must adhere to the use-based restrictions stated by the licensor, and these restrictions must also apply to derivative works. With a multitude of different model licenses available, it becomes a challenging and tedious work to reuse them in bulk. It is, therefore, imperative to establish guidelines for selecting the licenses for models and other related components that are ready for query-based FL.
IV-C2 License Choosing
@@
  • A model license ready for open FL paltforms should allow the modification, combination and redistribution of original works and any derived works.
@@
  • Sublicensing right should be granted to enable the republication of derived works after knowledge mining.
@@
  • Some licenses enforce the source of the derived works to be disclosed and prohibit their commercial use, which hinders model selling [109].
@@
  • Some licenses are copyleft (marked with * in TABLE II), which means the derivatives must be licensed under the same license or a compatible license, leading to potential license conflicts and proliferation [110].
  • All granted rights are preferably irrevocable by the licensors [111].
@@
Furthermore, it is important to consider the licensing of two other components when building and reusing models: data and algorithms, which may have entirely different license terms. Here, guided by the comparisons between different licenses outlined in TABLE II, we can summarize the preferences for selecting licenses in query-based FL to minimize conflicts‡III. Once we obtain the right to relicense the modification models, the choice of a new license depends on the application scenario of models. We further provide a flowchart in Fig. 5(a) to guide the license selection in the context of model query and model reusing. In addition, to enhance the copyrightability of the reused models, we should avoid using any derivatives and generated content from models under proprietary licenses throughout the ML reusing lifecycle‡III-D.
@@
V Batch Model Reuse Mechanisms
@@
In this section, we delve into the typical batch model reuse mechanisms from a technical perspective and introduce a new taxonomy designed to address the mismatch between license terms and technical terms. Therefore, instead of summarizing the batch model reuse mechanisms from a technological and algorithmic aspect, we propose grouping these mechanisms based on the classification of their resulting outputs for ease of justifying license clauses. As shown in Fig. 5(b), there are four categories of batch model reuse mechanisms: Combination, Amalgamation, Distillation, and Generation, each resulting in different forms of outputs and may correspond to different regulations in licenses.
@@
Figure 5: (a) Flowchart for model licenses selection in the context of model query and model reusing. (b) Proposed taxonomy categorizing batch model reuse mechanisms based on the reused results.
V-A Combination
@@
Combination [81] is a straightforward way to reuse batch of models (base learners), in which multiple models jointly contribute to the output by combination strategies such as averaging, voting, stacking [76, 112]. For regression estimates, averaging can improve the generalization by taking the mean of the outputs of all weak learners in a population. Additionally, the outputs of each learner can be weighted by extra parameters [113], which can be determined by stacking estimators [112], Bayesian approach [114] or backpropagation of gating networks [76]. Voting is a workaround strategy for classification tasks and also applicable for stacking and gating. Both stacking and gating rely on a holdout or validation dataset for calculating extra parameters, marked as * in Fig. 5(b). The difference is that gating can adapt the weights of each model’s estimation based on the inputs, providing better generalizability performance of the combined model.
@@
There are many advantages of combination mechanisms from the perspective of FL. First, the input spaces of base models can be unaligned, which is ideal for the scenario of vertical FL [115] where each client may have inconsistent features in their data. Secondly, especially for query-based FL, it can simultaneously support multiple types and heterogeneous models, which means that it does not rely on any prior assumptions of the models, such as whether they are DNNs or decision trees, released with raw weights (whitebox) or binary forms (blackbox). Thirdly, the tasks of models can be different if we pipeline the base models end-to-end, which is usually overlooked as a combination mechanism of models. Pipelining can fully leverage the transferability of models to solve previously unexplored ML problems. For instance, Gao et al. [116] proposed a zero-shot dense retrieval system named HyDE by pipelining a natural language generation (NLG) model [117] and a natural language understanding (NLU) model [118]. The generated content, which may lack factual grounding, from the NLG model is used as query embeddings to facilitate real document retrieval by the NLU model. Similarly, through query-based FL, we can query a vicarious NLG model for a novel scenario, such as ProGen [119] for protein sequences generation, and quickly adapt this system to proteomics. Not limited to that, we can query a batch of NLG models by a well-chosen filter condition and then combine models through averaging or gating to significantly expand the exploration space for knowledge discovery.
Lastly, the combinated models have strong separation from each other, meaning that we can add or remove a batch of models without significant changes to the remaining ones. Meanwhile, combination mechanisms do not rely on the transparency of models and support blackbox sharing. Thus, the base model can establish loose connections with other models only through run scripts, providing revocability of such combination and circumvention of the restrictions of licenses. On the other hand, instead of being treated as a challenge for FL [28], the statistical heterogeneity and model heterogeneity nature of these crowdsourced models can actually enrich population diversity, which is crucial for creating a good ensemble [120, 121].
@@
However, the storage consumption of combined models increases linearly with the number of base learners, which can strain the communication resources of a collaborative model training network. As a result, alternative approaches involve amalgamating or merging multiple models to create a new consensus model. In the following, we provide a summary of these methods.
V-B Amalgamation